OpenStudy (anonymous):

Which situation does the graph represent? A. Tom starts a rock collection. Each week he adds 15 more rocks. The graph shows how many rocks Tom has each week. B. Tom has 5 rocks in his collection. Each week he adds 15 more rocks. The graph shows how many rocks Tom has each week. C. Tom has 10 rocks in his collection. Each week he adds 5 more rocks. The graph shows how many rocks Tom has over time. D. Tom has 15 rocks in his collection. Each week he adds 5 more rocks. The graph shows how many rocks Tom has over time.

OpenStudy (anonymous):

@mLe

OpenStudy (anonymous):

I think the answer is C. I am just not 100%

OpenStudy (mle):

when weeks = 0 how many rocks does he have? aka what value does the graph start with?

OpenStudy (anonymous):

it starts with 15 (0,15)

OpenStudy (mle):

yes, so which option do you think it should be now?

OpenStudy (anonymous):

next point is (10,65)

OpenStudy (mle):

but if he Starts with 15 and *then* adds more rock to the 15 he starts with over time...

OpenStudy (anonymous):

if in ten weeks he has 65 and he started with 15 then 65-15=50; 50/5=10....that is why i think c

OpenStudy (mle):

65-15 = 50 10 - 0 = 10 50/10 = 5 he adds 5 rocks to his collection each week

OpenStudy (mle):

starts with 15 rocks them add 5 rocks a week

OpenStudy (anonymous):

so i am right yay!

OpenStudy (anonymous):

i srry... i meant D. is what i put...lol

OpenStudy (anonymous):

i'm a brain dead monkey head today :)

OpenStudy (mle):

ohhhhh ok than xD

OpenStudy (anonymous):

can u help with another graph? i don't get it bc there doesn't seem to b enough info... not to me anyway...lol